OpenStudy (paxpolaris):

How do I integrate cos^2 ? \[\large\int_0^{2\pi} \cos^2u\ du\]

OpenStudy (helder_edwin):

use this identity \[ \large \cos^2u=\frac{1+\cos(2u)}{2} \]

OpenStudy (paxpolaris):

d'oh ... Thanks ... u r awesome